OpenStudy (anonymous):

use the distributive property to simplify the expression. 2x-3 [ 5x-2(6-x)] A. 8x-11 B. 19x -36 C. -19x + 36 D. -8+11

OpenStudy (anonymous):

36 - 19 x

OpenStudy (anonymous):

2x-3 [ 5x-2(6-x)] =2x-3[5x-12+2x]=2x-3[7x-12]=2x-21x+36=-19x+36

OpenStudy (anonymous):

i appriciate you guys helping em with my final